# Learning Constraints from Stochastic Partially-Observed Closed-Loop Demonstrations

### v1

@@ III-C Sensitivity Analysis With Respect to Transmission Error
We now prove that when the transmission noise $\{(\bm{\delta}_{u}^{({d})},\bm{\delta}_{y}^{({d})}):d\in[D]\}$ is nonzero, the error in the recovered output feedback $\mathcal{K}$ and nominal trajectory $(\textbf{z},\textbf{v})$ scale at most linearly in the noise magnitudes $\|\bm{\delta}_{u}^{({d})}\|_{2}$ and $\|\bm{\delta}_{y}^{({d})}\|_{2}$.
Theorem 3
@@
Suppose Assumption 1 holds, and $\max_{d\in[D]}\max\{\|\bm{\delta}_{u}^{({d})}\|_{2},\|\bm{\delta}_{y}^{({d})}\|_{2}\}<\epsilon$ for some $\epsilon>0$. Define $\textbf{Y}:=\begin{bmatrix}\textbf{y}^{({2})}-\textbf{y}^{({1})}&\cdots&\textbf{y}^{({D})}-\textbf{y}^{({D-1})}\end{bmatrix}$, $\rho_{1}=\sqrt{D-1}\|\textbf{Y}^{\dagger}\|_{2}$, $\rho_{2}=\sqrt{D-1}\|\Gamma\|_{2}\|\mathcal{C}\|_{2}$, $\rho_{3}=\sqrt{D-1}\|\Gamma\|_{2}\|\textbf{Y}^{\dagger}\|_{2}$, and $\rho_{4}=\|\Gamma\|_{2}\|\mathcal{K}^{\star}\|_{2}.$ Let $(\delta\textbf{z},\delta\textbf{v},\delta K):=(\tilde{\textbf{z}}-\textbf{z}^{\star},\tilde{\textbf{v}}-\textbf{v}^{\star},\tilde{\mathcal{K}}-\mathcal{K}^{\star})$. Then:
@@
$\left\|\begin{bmatrix}\delta\textbf{z}\\ \delta\textbf{v}\end{bmatrix}\right\|_{2}$ $\leq\left(\rho_{2}\left\|\begin{bmatrix}\tilde{\textbf{z}}\\ \tilde{\textbf{v}}\end{bmatrix}\right\|_{2}+\rho_{3}\|\tilde{\textbf{y}}^{({1})}\|_{2}(\|\tilde{\mathcal{K}}\|_{2}+1)+\rho_{4}\right)\epsilon.$
Proof:
@@
IV Experiments
To evaluate our method, we perform constraint learning on simulated 4D double integrator, 4D unicycle, nonlinear 6D quadcopter, and linearized 12D quadcopter dynamics [19].
IV-A Experiment Setup
@@
$C_{\urcorner{k}}(\theta)$ $=\bigwedge_{\alpha\in[\overline{N}_{c}]}\bigvee_{\beta\in[N_{c}]}\big\{A_{\alpha,\beta}(\theta)(\textbf{x},\textbf{u})\leq b_{\alpha,\beta}(\theta)\big\},$
where $A_{\alpha,\beta}(\theta)$ and $b_{\alpha,\beta}$ are dimension-compatible $\forall\alpha,\beta$.
@@
In the Figs. 2c, 3a-b, 4a-c simulations, the set $W$ bounding the dynamics noise terms $w_{t}$ was defined to be $\infty$-norm balls of radii 0.05, 0.05, 0.05, 0.2, 0.07, and 0.05, respectively; in all other simulations, $W$ was set as the unit $\infty$-norm ball. In the Fig. 2 simulations, the set $V$ bounding the dynamics noise terms $v_{t}$ was set as the unit norm ball of radius 0.02; all other simulations used zero output noise.
@@
As shown in Figs. 2-4, for each experiment, we fix a ground truth constraint set (shaded gray, with yellow boundaries), and use output SLS, control contraction metrics (CCM), or proportional–derivative (PD) control to generate nominal trajectories (green, with circles/stars for origins/goals), output feedback laws, and demonstration rollouts (blue) with corresponding reachable tubes and tube edges (magenta). We then apply the methods in Sec. III to learn the nominal trajectories, output feedback laws, and constraints (boundaries in black) from the given demonstrations.
For each experiment, we generate demonstrations either via (full) state feedback (i.e., $y_{t}=x_{t}$), or output feedback, wherein only the position vector is observed (i.e., $y_{t}=p_{t}$). For the baseline experiment (Fig. 2b), we use $D=20$ demonstrations; for all others, we set $D=100$. The simulations in Figs. 2-4 are generated with zero transmission noise; Fig. 1 shows that the learning accuracy of our method degrades gracefully as transmission error magnitude increases.
@@
Figure 1: Error magnitude in the recovered nominal trajectories, nominal controls, output feedback, and constraint parameter as a function of the transmission error level. Demonstrations were generated using noise-corrupted double integrator dynamics and (full) state feedback SLS controllers. On average, our method enjoys high accuracy in learning a priori unknown parameters under low levels of transmission noise.
IV-B Simulation Results
@@
Figure 2: Constraint learning from demonstrations generated via SLS with double integrator (a, c) and linear 12D quadcopter (d, e, f) dynamics. (e, f) provide views of (d) from different angles. (a, c-f) Our method accurately learns the true collision avoidance constraints across all simulations. (b) In contrast, the baseline [1] failed to accurately recover the true constraint; note the mismatch between the learned (black) and true (yellow) constraints.
SLS Controllers
@@
We present simulation results evaluating our constraint recovery method (Figs. 2a, c-f) and comparing it to the baseline in [1] (Fig. 2b), which was designed to learn constraints from noise-free, fully observed state trajectories. We learn constraints from demonstrations generated by unrolling noise-corrupted double integrator (Fig. 2 a-c) or linearized 12D quadcopter (Fig. 2d-f) dynamics, and applying a state feedback (Fig. 2 a-b, d-f) or output feedback (Fig. 2c) SLS controller (Fig. 2c). The nominal trajectories and feedback laws underlying the input-output demonstration were jointly generated via (14). Across experiments in Figs. 2a, c-f, our methods, as described in (23)-(26), accurately learned the true nominal trajectories, output feedback laws, and constraint parameters underlying the input-output demonstrations. In contrast, the baseline [1] fails to learn the true underlying constraints (Fig. 2b).
@@
Figure 3: Constraint learning from demonstrations produced via CCM with unicycle (a) and double integrator (b) dynamics. Our method accurately learns the true constraints for all simulations.
Control Contraction Metrics (CCM)-Based State Feedback Controllers
@@
We learn constraints from demonstrations generated by unrolling noise-corrupted unicycle (Fig. 3a) or double integrator (Fig. 3b) dynamics with full state observation, using the CCM-based state feedback law in [16]. Concretely, we computed the error feedback law $\mathcal{K}$ prescribed in [16, Sec. III-B], and the associated system response $\Phi$ via (11). We then use (14) to compute the nominal trajectory for generating demonstrations, while fixing $\Phi$ fixed at the value derived from via CCM (see Remark 2). Across experiments in Fig. 3, by applying (23)-(26), we accurately learned the true nominal trajectories, output feedback laws, and constraint parameters underlying the demonstrations.
@@
Figure 4: Constraint learning from demonstrations produced via CCM with (a) unicycle, (b) double integrator, and (c) nonlinear 6D quadcopter dynamics. Our method accurately learns the true constraints across all simulations.
Proportional-Derivative (PD) Controllers
@@
We recover constraints from demonstrations generated using noise-corrupted unicycle (Fig. 4a), double integrator (Fig. 4b), or nonlinear 6D quadcopter (Fig. 4c) dynamics, by applying PD state feedback controllers. Similar to the CCM-based methods, we first compute the feedback policy $\mathcal{K}$ and the corresponding system response $\Phi$ via PD control techniques and (11). To obtain a nominal trajectory from which demonstrations are then generated, we solve (14) while holding $\Phi$ fixed at its value computed via the PD control method and (11). In the Fig. 4 experiments, our methods, as described by (23)-(26) accurately recovered the true nominal trajectories, output feedback laws, and constraint parameters that characterize the input-output demonstrations.
V Conclusion

### v2

@@ 3.3 Sensitivity Analysis With Respect to Transmission Error
We now prove that when the transmission noise $\{(\bm{\delta}_{u}^{({d})},\bm{\delta}_{y}^{({d})}):d\in[D]\}$ is nonzero, the error in the recovered output feedback $\mathcal{K}$ and nominal trajectory $(\textbf{z},\textbf{v})$ scales at most linearly in the noise magnitudes $\|\bm{\delta}_{u}^{({d})}\|_{2}$ and $\|\bm{\delta}_{y}^{({d})}\|_{2}$.
Theorem 3.7.
@@
Suppose $f_{t}(\cdot)$ is LTV, Assumption 1 holds, and $\max_{d\in[D]}\max\{\|\bm{\delta}_{u}^{({d})}\|_{2},\|\bm{\delta}_{y}^{({d})}\|_{2}\}<\epsilon$ for some $\epsilon>0$. Suppose $\textbf{Y}:=\begin{bmatrix}\textbf{y}^{({2})}-\textbf{y}^{({1})}&\cdots&\textbf{y}^{({D})}-\textbf{y}^{({D-1})}\end{bmatrix}$ has full row rank, and set $\rho_{1}:=\sqrt{D-1}\|\textbf{Y}^{\dagger}\|_{2}$, $\rho_{2}:=\sqrt{D-1}\|\textbf{Y}^{\dagger}\|_{2}\|\Gamma^{-1}\|_{2}\|\mathcal{C}\|_{2}$, $\rho_{3}:=\sqrt{D-1}\|\textbf{Y}^{\dagger}\|_{2}\|\Gamma^{-1}\|_{2}$, and $\rho_{4}:=\|\Gamma^{-1}\|_{2}\big(\|\mathcal{K}^{\star}\|_{2}+1\big).$ Let $(\delta\textbf{z},\delta\textbf{v},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\delta\mathcal{K}}):=(\tilde{\textbf{z}}-\textbf{z}^{\star},\tilde{\textbf{v}}-\textbf{v}^{\star},\tilde{\mathcal{K}}-\mathcal{K}^{\star})$. Then:
@@
$\left\|\begin{bmatrix}\delta\textbf{z}\\ \delta\textbf{v}\end{bmatrix}\right\|_{2}$ $\leq\left(\rho_{2}\left\|\begin{bmatrix}\tilde{\textbf{z}}\\ \tilde{\textbf{v}}\end{bmatrix}\right\|_{2}+\rho_{3}\|\tilde{\textbf{y}}^{({1})}\|_{2}\right)(\|\tilde{\mathcal{K}}\|_{2}+1)\epsilon+\rho_{4}\epsilon.$
Proof 3.8.
@@
4 Experiments
To evaluate our method, we perform constraint learning on simulated 4D double integrator, 4D unicycle, nonlinear 6D quadcopter, and linearized 12D quadcopter dynamics [15]. We use Gurobi to solve all optimization problems. Our code and additional results can be found at https://github.com/zhangzdd/SLS-ConstraintLearning.
4.1 Experiment Setup
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}S_{\urcorner{k}}}(\theta)$ $=\wedge_{\alpha\in[\overline{N}_{c}]}\vee_{\beta\in[N_{c}]}\big\{A_{\alpha,\beta}(\theta)(\textbf{x},\textbf{u})\leq b_{\alpha,\beta}(\theta)\big\},$
where $A_{\alpha,\beta}(\theta)$ and $b_{\alpha,\beta}(\theta)$ are dimension-compatible $\forall\alpha,\beta$.
@@
In simulations for Figs. 2c, 3a, 3b, 4a, 4b, 4c, the set $W$ bounding the terms $w_{t}$ was respectively defined to be $\infty$-norm balls of radii 0.05, 0.05, 0.05, 0.2, 0.07, and 0.05; in all other simulations, $W$ was set as the unit $\infty$-norm ball. For Fig. 2, the set $V$ bounding the output noise terms $e_{t}$ was set as the norm ball of radius 0.02; all other simulations used zero output noise. As shown in Figs. 2-4, for each experiment, we fix a ground truth constraint set (shaded gray), and use output SLS, control contraction metrics (CCM), or proportional–derivative (PD) control to generate nominal trajectories (green, with circles/stars for origins/goals), output feedback laws, and demonstration rollouts (blue) with corresponding state trajectory bounds (edges in magenta) [^4] Such bounds are computed using $\Phi(\textbf{w},\textbf{e})$, the maximum deviation from the nominal, and realized closed-loop, trajectories (Sec. 2.2).. We then apply our methods (Sec. 3) to learn nominal trajectories, output feedback laws, and constraints (boundaries in black) from the given demonstrations. For each experiment, we generate demonstrations either via (full) state feedback (i.e., $y_{t}=x_{t}$), or output feedback, wherein only the position vector is observed (i.e., $y_{t}=p_{t}$), and we set $D=100$. The simulations in Figs. 2-4 are generated with zero transmission noise; Fig. 1 shows that the learning accuracy of our method degrades gracefully as transmission error magnitude increases.
@@
Figure 1: Error magnitude in recovered (a) nominal trajectories and controls, (b) output feedback, and (c) constraint parameter as a function of the transmission error level. Demonstrations were generated using noise-corrupted double integrator dynamics and (full) state feedback SLS controllers. On average, our method learns a priori unknown parameters with high accuracy under low levels of transmission noise.
4.2 Simulation Results
@@
Figure 2: Constraint learning from demonstrations generated via SLS with double integrator (a, c) and linear 12D quadcopter (d, e, f) dynamics. (e, f) provide views of (d) from different angles. (a, c-f) Our method accurately learns true collision avoidance constraints across all simulations, but (b) the baseline [4] did not; note the mismatch between the learned (black) and true (yellow) constraints.
@@
a) SLS Controllers: We present simulations evaluating our method (Figs. 2a, c-f) and comparing it to the baseline in [4] (Fig. 2b), which was designed to learn constraints from noise-free, fully observed state trajectories. We learn constraints from demonstrations generated by unrolling noise-corrupted double integrator (Fig. 2 a-c) or linearized 12D quadcopter (Fig. 2d-f) dynamics, and applying a state feedback (Fig. 2 a-b, d-f) or output feedback (Fig. 2c) SLS controller (Fig. 2c). The nominal trajectories and feedback laws underlying the demonstration were jointly generated via (14). Across experiments in Figs. 2a, c-f, our methods, as described in (23)-(26), accurately learned the true nominal trajectories, output feedback laws, and constraint parameters underlying the input-output demonstrations. In contrast, the baseline [4] fails to learn the true underlying constraints (Fig. 2b).
@@
Figure 3: Constraint learning from demonstrations produced via CCM with unicycle (a) and double integrator (b) dynamics. Our method accurately learns true collision avoidance constraints across all simulations.
@@
b) Control Contraction Metrics (CCM)-Based State Feedback Controllers: We learn constraints from demonstrations generated by unrolling noise-corrupted unicycle (Fig. 3a) or double integrator (Fig. 3b) dynamics with full state observation, using the CCM-based state feedback law in [10]. Concretely, we computed the error feedback law $\mathcal{K}$ of [10, Sec. III-B], and the associated system response $\Phi$ via (11). We then use (14) to compute the nominal trajectory for generating demonstrations, while fixing $\Phi$ fixed at the value derived from via CCM (see Remark 2). By applying (23)-(26), we accurately learned the true nominal trajectories, output feedback laws, and constraint parameters (Fig. 3).
@@
Figure 4: Constraint learning from demonstrations produced via PD with (a) unicycle, (b) double integrator, and (c) nonlinear 6D quadcopter dynamics. Our method accurately learns true collision avoidance constraints for all simulations.
@@
c) Proportional-Derivative (PD) Controllers: We recover constraints from demonstrations generated using noise-corrupted unicycle (Fig. 4a), double integrator (Fig. 4b), or nonlinear 6D quadcopter (Fig. 4c) dynamics, by applying PD state feedback controllers. Similar to the CCM-based methods, we first compute the feedback policy $\mathcal{K}$ and the corresponding system response $\Phi$ via PD control techniques and (11). To obtain a nominal trajectory from which demonstrations are then generated, we solve (14) while holding $\Phi$ fixed at its value computed via the PD control method and (11). In the Fig. 4 experiments, our methods, as described by (23)-(26) accurately recovered the true nominal trajectories, output feedback laws, and constraint parameters that characterize the input-output demonstrations.
5 Conclusion